Using these data from the comparative balance sheet of Blossom Company, perform horizontal analysis. (If amount and percentage are a decrease show the numbers as negative, e.g. -55,000, -20% or (55,000), (20%). Round percentages to 0 decimal places, e.g. 12%.)

Increase or (Decrease)

Dec. 31, 2022

Dec. 31, 2021

Amount

Percentage

Accounts receivable

$ 345,000 $ 300,000 $Enter increase or decrease amount in dollars.

Enter increase or decrease in percentages rounded to 0 decimal places.

%

Inventory

$ 649,000 $ 550,000 $Enter increase or decrease amount in dollars.

Enter increase or decrease in percentages rounded to 0 decimal points.

%

Total assets

$3,024,000 $2,700,000 $Enter increase or decrease amount in dollars.

Enter increase or decrease in percentages rounded to 0 decimal points.

%
Answer #1

Increase or (Decrease)

Dec. 31, 2022

Dec. 31, 2021

Amount

Percentage

Accounts receivable

$ 345,000 $ 300,000 345,000 - 300,000 = 45,000

45,000/300,000 = 15

%

Inventory

$ 649,000 $ 550,000

649,000 - 550,000 = 99,000

99,000/550,000 = 18 %

Total assets

$3,024,000 $2,700,000

3,024,000 - 2,700,000 = 324,000

324,000/2,700,000 = 12

%